# SDE-SQL: Enhancing Text-to-SQL Generation in Large Language Models via Self-Driven Exploration with SQL Probes

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly improved performance on the Text-to-SQL task. However, prior approaches typically rely on static, pre-processed database information provided at inference time, which limits the model’s ability to fully understand the database contents. Without dynamic interaction, LLMs are constrained to fixed, human-provided context and cannot autonomously explore the underlying data. To address this limitation, we propose SDE-SQL, a framework that enables large language models to perform Self-Driven Exploration of databases during inference. This is accomplished by generating and executing SQL probes, which allow the model to actively retrieve information from the database and iteratively update its understanding of the data. Unlike prior methods, SDE-SQL operates in a zero-shot setting, without relying on any question-SQL pairs as in-context demonstrations. When evaluated on the BIRD benchmark with Qwen2.5-72B-Instruct, SDE-SQL achieves an 8.02% relative improvement in execution accuracy over the vanilla Qwen2.5-72B-Instruct baseline, establishing a new state-of-the-art among methods based on open-source models without supervised fine-tuning (SFT) or model ensembling. Moreover, with SFT, the performance of SDE-SQL can be further enhanced, yielding an additional 0.52% improvement.
@@
Wenxuan Xie1, Yaxun Dai2, Wenhao Jiang3thanks: Corresponding author 1South China University of Technology  2Soochow University  3Guangdong Laboratory of AI and Digital Economy (SZ) lancelotxie601@gmail.com, cswhjiang@gmail.com
@@ 1 Introduction
Text-to-SQL is a long-standing research task in the field of natural language processing, aiming to translate natural language questions into SQL queries that can be executed by a database. It not only enables non-experts to interact with databases effortlessly, but also helps alleviate hallucinations in question-answering systems by providing LLMs with knowledge stored in the database.
Recent developments in LLMs have substantially boosted the performance and accuracy of Text-to-SQL applications. Methods based on LLMs have already achieved over 90% execution accuracy on the early Spider dataset (Yu et al., 2019), and have also shown good performance on the more complex and challenging BIRD dataset. However, a gap still exists between these methods and human experts, which is particularly evident in the recently proposed Spider 2.0 benchmark (Lei et al., 2025).
@@
Current mainstream large-model-based Text-to-SQL methods typically consist of three key modules: schema linking, SQL generation, and SQL refinement. In the schema linking stage, previous work has primarily focused on matching more precise and relevant database information to specific natural language questions. In the SQL generation stage, various question decomposition methods and reasoning strategies have been proposed. In the SQL refinement stage, error types have been more thoroughly categorized, leading to the development of more targeted correction techniques. However, one critical yet often overlooked aspect of SQL is its role as an interactive interface with the database, enabling rapid and efficient execution. This property remains underutilized in current approaches and may partly explain the performance gap between LLM-based systems and human experts.
Therefore, we propose SDE-SQL, introducing the Self-Driven Exploration technique in both the generation and refinement stages, which is shown in Figure 1. In addition to the final SQL query that corresponds to the natural language question, a series of SQL queries specifically designed for database exploration are generated and executed during the task. These SQL queries are referred to as SQL Probes. For schema linking, we employ entity-based value retrieval and entity-based soft linking methods. In the generation stage, we allow the LLMs to independently explore the database in two phases based on the current natural language question and the database schema, and then enable LLMs to perform zero-shot learning with the exploration results.
After generation, we refine incorrect SQL queries using two strategies. For those returning explicit error messages, LLMs directly revise the query based on the feedback. For SQLs that return empty results, we apply rule-based decomposition to generate a series of sub-queries (Sub-SQLs). LLMs then analyze their execution results and perform targeted exploration to identify and fix the underlying issues.
When evaluated on the BIRD dataset, SDE-SQL based on Qwen2.5-72B-Instruct achieved an execution accuracy of 67.67%. After SFT, the performance further improves to 68.19%.
In summary, here are our contributions:
  • We propose SDE-SQL, a novel framework that leverages Self-Driven Exploration to unlock the true potential of LLMs in the Text-to-SQL task, allowing them to use SQLs as probes to explore the database and thereby reducing the performance gap between LLMs and human experts.
  • We introduce a two-phase exploration in the SQL generation stage, empowering LLMs with the ability to dynamically interact with the database, thereby breaking the limitations of information availability.
  • In the SQL refinement stage, we use a rule-based approach to decompose the SQL into a series of Sub-SQLs, enabling LLMs to analyze the execution results of these Sub-SQLs, diagnose the error sources with the execution results of generated SQL Probes, and further interact with the database to explore solutions based on the diagnostic conclusion.
  • We evaluated our framework on the BIRD dataset and Spider dataset and conducted a series of ablation experiments, demonstrating the importance and effectiveness of Self-Driven Exploration for the Text-to-SQL task.
  • We also constructed a small amount of data for Supervised Fine-Tuning (SFT) on the exploration and generation tasks separately. The results show that fine-tuning individual modules can lead to improvements in the overall performance of SDE-SQL.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have led to substantial progress on the Text-to-SQL task. However, existing approaches typically depend on static, pre-processed database information supplied at inference time, which restricts the model’s capacity to deeply comprehend the underlying database content. In the absence of dynamic interaction, LLMs are limited to fixed, human-curated context and lack the ability to autonomously query or explore the data. To overcome this limitation, we introduce SDE-SQL, a novel framework that empowers LLMs to perform Self-Driven Exploration of databases during inference. This is achieved through the generation and execution of SQL probes, enabling the model to actively retrieve information and iteratively refine its understanding of the database. Unlike prior methods, SDE-SQL operates in a zero-shot setting, requiring no in-context demonstrations or question-SQL pairs. Evaluated on the BIRD benchmark with Qwen2.5-72B-Instruct, SDE-SQL achieves an 8.02% relative improvement in execution accuracy over the vanilla Qwen2.5-72B-Instruct baseline, establishing a new state-of-the-art among open-source methods without supervised fine-tuning (SFT) or model ensembling. Furthermore, when combined with SFT, SDE-SQL delivers an additional 0.52% performance gain.
@@
Wenxuan Xie1, Yaxun Dai2, Wenhao Jiang3thanks: Corresponding author 1South China University of Technology, 2Soochow University 3Guangdong Laboratory of AI and Digital Economy (SZ) lancelotxie601@gmail.com, cswhjiang@gmail.com
@@ 1 Introduction
Text-to-SQL is a long-standing task in natural language processing that focuses on translating natural language questions into executable SQL queries. This capability not only empowers non-expert users to interact with structured databases seamlessly, but also mitigates hallucination issues in question-answering systems by grounding responses in factual, database-stored information.
Recent advances in large language models (LLMs) have led to significant improvements in the performance and accuracy of Text-to-SQL systems. LLM-based approaches have surpassed 90% execution accuracy on the original Spider dataset (Yu et al., 2019), and have demonstrated promising results on more complex and diverse benchmarks such as BIRDLi et al. (2023). Despite these advances, a noticeable gap remains between current model performance and human-level capabilities—particularly on the recently introduced Spider 2.0 benchmark (Lei et al., 2025), which poses more realistic and challenging scenarios for semantic parsing.
@@
Contemporary large language model (LLM)-based approaches to Text-to-SQL typically comprise three core components: schema linking, SQL generation, and SQL refinement. In the schema linking stage, prior work has primarily focused on aligning natural language questions with relevant database schema elements, improving precision and contextual relevance. During SQL generation, various methods have been proposed to decompose complex questions and incorporate reasoning strategies. In the refinement stage, the categorization of SQL error types has become more systematic, enabling the development of targeted correction mechanisms.
Despite these advances, one crucial aspect of SQL remains largely underexplored: its inherent interactivity as a database interface that supports fast and informative execution. This underutilized property may partially account for the performance gap between LLM-based systems and human experts.
To address this, we propose SDE-SQL, a novel framework that incorporates Self-Driven Exploration into both the generation and refinement stages, as illustrated in Figure 1. In addition to generating the final SQL query that directly answers the natural language question, the model autonomously generates and executes a sequence of auxiliary queries—termed SQL Probes—designed specifically to explore and extract informative signals from the database.
For schema linking, we leverage entity-based techniques including value retrieval and soft linking. During the generation phase, the model engages in a two-stage exploration process based on the question and schema, enabling it to iteratively refine its understanding of the database content and perform zero-shot reasoning grounded in the retrieved information.
Following generation, we incorporate a two-stage exploration process into the refinement phase. For SQL queries that return explicit execution errors, the model directly revises them based on the error feedback. For queries that execute successfully but return empty results, the first stage of exploration uses the execution results of decomposed sub-queries (Sub-SQLs) to help the model diagnose the underlying issue. In the second stage, the model generates targeted SQL Probes to explore possible solutions, and selects the most promising one to produce the final refined query.
Empirically, SDE-SQL achieves an execution accuracy of 67.67% on the BIRD benchmark using Qwen2.5-72B-Instruct in a zero-shot setting. With supervised fine-tuning (SFT), the performance further improves to 68.19%.
Our main contributions are as follows:
  • We propose SDE-SQL, a novel framework that leverages Self-Driven Exploration to enhance the reasoning and interaction capabilities of LLMs in the Text-to-SQL task, significantly narrowing the gap with human experts.
  • We introduce a unified exploration mechanism across both SQL generation and refinement stages, enabling LLMs to actively query the database, diagnose potential errors, and iteratively improve query quality.
  • We conduct extensive experiments on the BIRD and Spider benchmarks, along with ablation studies, validating the effectiveness of Self-Driven Exploration.
  • We build a small-scale dataset for supervised fine-tuning (SFT) on exploration and generation tasks, and show that targeted module-level fine-tuning further improves the performance of SDE-SQL.